# Jacquet modules of Tate cohomology and base change lifting

### v1

@@ 2. Tate cohomology of smooth representations
Let $\Pi$ be a smooth finite length representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$, we show that the Tate cohomology groups $\widehat{H}^{i}(\Gamma,\Pi)$ are finitely generated representations of $G^{\sigma}(F)$. We will use Schneider–Stuhler resolutions and their $-\ell$ version by Vignéras for proving this finiteness statement.
@@
We fix an algebraic closure $\overline{\mathbb{Q}}_{\ell}$ of $\mathbb{Q}_{\ell}$ and let $\mathcal{K}$ be the maximal unramified extension of $\mathbb{Q}_{\ell}$ in $\overline{\mathbb{Q}}_{\ell}$. Let $\Lambda$ be the ring of integers of $\mathcal{K}$. All $\Lambda[G(F)]$ modules are assumed to be smooth, i.e., every element is fixed by an open subgroup of $G(F)$. Let $M$ be a $\Lambda[G(F)\rtimes\Gamma]$ module. Let ${\rm Nr}$ be the norm element $\sum_{g\in\Gamma}g\in\Lambda[\Gamma]$. Then the cohomology of the following chain complex of $\Lambda[\Gamma]$ modules
@@
is $2$-periodic. The Tate cohomology groups of $M$ are defined as
@@
The action of $G(F)$ on $M$ induces an action of $G^{\sigma}(F)$ on each Tate cohomology group $\widehat{H}^{i}(\Gamma,M)$ and we get two $\ell$-modular smooth representations $\widehat{H}^{i}(\Gamma,M)$, for $i\in\{0,1\}$, of the fixed point subgroup $G^{\sigma}(F)$. Similar cohomology groups can be defined when $F$ is a finite field, and we will assume the easy modifications in the definitions.
@@
Let $X$ be a Hausdorff, locally compact and totally disconnected topological space (an $\ell$ space in the sense of [BZ76]). Let $G$ be a locally profinite group with an order $\ell$ automorphism $\sigma$, generating a group $\Gamma$ of automorphisms of $G$. Assume that $G\rtimes\Gamma$ acts continously on $X$. Let $\mathcal{F}$ be a $G\rtimes\Gamma$ equivariant sheaf on $X$. The natural restriction map
@@
is an isomorphism (see [TV16, Proposition 3.3]). Let $H$ be a closed subgroup of $G$ such that $\sigma(H)=H$. Let $V$ be a $\Lambda[H\rtimes\Gamma]$ module. This data defines a $G\rtimes\Gamma$ equivariant sheaf $\mathcal{F}_{V}$ on $G/H$ whose sections on an open set $U$ of $G/H$ are the sections of the vector bundle $G\times_{H}V\rightarrow G/H$ restricted to $U$. The global sections of the sheaf $\mathcal{F}_{V}$ is the induced representation ${\rm ind}_{H}^{G}V$, and the action of $\Gamma$ is given by
$(\sigma.f)(g):=\sigma\big{(}f(\sigma^{-1}(g))\big{)},$
for all $g\in G$ and $f\in{\rm ind}_{H}^{G}V$. Now, we consider the Tate cohomology groups $\widehat{H}^{i}(\Gamma,{\rm ind}_{H}^{G}V)$, $i\in\{0,1\}$. Assume that $(G/H)^{\sigma}=G^{\sigma}/H^{\sigma}$. Then, the restriction to $G^{\sigma}$ map induces a $G^{\sigma}$ equivariant isomorphism
@@
Let $(\Pi,V)$ be a finite length $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$-representation. We will show that $\widehat{H}^{i}(\Gamma,\Pi)$ is a finitely generated $G^{\sigma}(F)$ representation. We will the mod $\ell$ version of Schneider and Stuhler resolutions ([SS97, Theorem II.3.1]) of smooth irreducible representations proved by Vignéras ([Vig97, Proposition 2.6]). Let $\mathfrak{X}$ be the extended Bruhat–Tits building of a connected reductive group $G$ defined over a non-Archimedean local field $F$. Let $\mathfrak{X}_{d}$ be the set of $q$-facets of $\mathfrak{X}$, and let $\mathfrak{X}_{(d)}$ be the set of oriented $q$-facets which are tuples of the form $(\mathcal{F},c_{\mathcal{F}})$ where $\mathcal{F}\in\mathfrak{X}_{d}$ and $c_{\mathcal{F}}$ is an orientation of $\mathcal{F}$. Let $e\geq 0$. For a facet $\mathcal{F}$ of $\mathfrak{X}$, Schneider and Stuhler associate a compact open subgroup $U_{\mathcal{F}}^{(e)}$, and let $\underline{V}$ be the coefficient system defined by associating every facet $\mathcal{F}$ to $V^{U^{e}_{\mathcal{F}}}$. Let $C(\mathfrak{X}_{(d)},\underline{V})$ be the set of oriented cellular $q$-chains on $\mathfrak{X}_{(d)}$. Then we have a $G$ equivariant resolution of the representation
$0\rightarrow C(\mathfrak{X}_{(d)},\underline{V})\xrightarrow{\partial_{d}}C(\mathfrak{X}_{(d-1)},\underline{V})\xrightarrow{\partial_{d-1}}\cdots\rightarrow C(\mathfrak{X}_{(1)},\underline{V})\xrightarrow{\partial_{1}}C(\mathfrak{X}_{(0)},\underline{V})\xrightarrow{\partial_{0}}V\rightarrow 0$ (2.2)
Note the automorphism $\sigma$ acts on $\mathfrak{X}$ simplicially, i.e., it preserves the sets $\mathfrak{X}_{d}$ and $\mathfrak{X}_{(d)}$.
Lemma 2.1.
The Tate cohomology groups $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d)},\underline{V}))$ are finitely generated $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$ modules.
@@
Consider the set $\mathfrak{X}_{(d)}$ with discrete topology, and the action of the group $G(F)$ on $\mathfrak{X}_{(d)}$ is continuous. Also note that the sheaf $\underline{V}$ is $G$ and $\langle\sigma\rangle$ equivariant. Since $V$ is an admissible representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$, the stalks of the sheaf $\underline{V}$ are all finite dimensional vector spaces over $\overline{\mathbb{F}}_{l}$. Using [TV16, Proposition 3.3], we get that the restriction map
$\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d)},\underline{V}))\rightarrow C(\mathfrak{X}_{(d)}^{\sigma},\widehat{H}^{i}(\Gamma,\underline{V}))$
is an isomorphism. Let $\mathcal{Y}$ be the Bruhat–Tits building of $G^{\sigma}$, then $\mathfrak{X}^{\sigma}_{d}$ is a subset of facets of $\mathcal{Y}$. Thus, there are only finitely many orbits for the action of $G^{\sigma}$ on $\mathfrak{X}^{\sigma}_{(d)}$. Since the module $\widehat{H}^{i}(\Gamma,V^{U_{F}^{e}})$ is finite, we get that the space $C(\mathcal{X}_{(d)},\widehat{H}^{i}(\Gamma,\underline{V}))$ is a finitely generated representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$. This shows the lemma. ∎
We will use the fact that the sub modules of finitely generated $\overline{\mathbb{F}}_{\ell}[H(F)]$ smooth modules are again finitely generated. This is a fundamental result in the theory of smooth complex representations due to Bernstein and in the $-\ell$ case it follows from the deep results of the paper [DHKM24, Corollary 1.3] seen with the work of Dat ([Dat09, Corollary 4.5].
Lemma 2.2.
Let $(\Pi,V)$ be an irreducible representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$. Assume that $(\Pi,V)$ extends as a representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. Then the Tate cohomology groups $\widehat{H}^{i}(\Gamma,V)$ are finitely generated representations of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$.
@@
Note that the resolution (2.2) is $G$ and $\Gamma$ equivariant. Tate cohomology groups
@@
are finitely generated representations of $\overline{\mathbb{F}}_{l}[G^{\sigma}(F)]$ by Lemma 2.1. For each $k$, consider the exact sequence $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$ modules
$0\rightarrow{\rm Im}(\partial_{k})\rightarrow C(\mathfrak{X}_{(k-1)},\underline{V})\rightarrow{\rm Ker}(\partial_{k-1})\rightarrow 0$
Then the corresponding long exact sequence of Tate cohomology gives the following hexagonal exact sequence
${\widehat{H}^{0}(\Gamma,{\rm Im}(\partial_{k}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{0}(\Gamma,C(\mathfrak{X}_{(k-1)},\underline{V}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{0}(\Gamma,{\rm Ker}(\partial_{k-1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{1}(\Gamma,{\rm Ker}(\partial_{k-1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{1}(\Gamma,C(\mathfrak{X}_{(k-1)},\underline{V}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{1}(\Gamma,{\rm Im}(\partial_{k}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
which we denote by $S(k)$. Note that ${\rm Im}(\partial_{d})$ is identified with $C(\mathfrak{X}_{d},\underline{V})$. Since both the Tate cohomology groups $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d)},\underline{V}))$ and $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d-1)},\underline{V}))$ are finitely generated, it follows from the diagram $S(d)$ that $\widehat{H}^{i}(\Gamma,{\rm Ker}(\partial_{d-1}))$ is also finitely generated. A similar hexagonal diagram of Tate cohomology corresponding to the following short exact sequence of $G(F)\rtimes\Gamma$ modules
$0\rightarrow{\rm Ker}(\partial_{d-1})\rightarrow C(\mathfrak{X}_{(d-1)},\underline{V})\rightarrow{\rm Im}(\partial_{d-1})\rightarrow 0$
shows that $\widehat{H}^{i}(\Gamma,{\rm Im}(\partial_{d-1}))$ is finitely generated. Since $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d-2)},\underline{V}))$ is finitely generated, it follows from the hexagonal diagram $S(d-1)$ that $\widehat{H}^{i}(\Gamma,{\rm Ker}(\partial_{d-2}))$ is finitely genereated. Thus, following a finite number of inductive steps, the hexagonal diagram $S(2)$ shows that the Tate cohomology $\widehat{H}^{i}(\Gamma,{\rm Ker}(\partial_{1}))$ is finitely generated. Finally, it follows from the hexagonal diagram of Tate cohomology groups corresponding to the following short exact sequence of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$ modules
$0\rightarrow{\rm Ker}(\partial_{1})\rightarrow C(\mathfrak{X}_{(0)},\underline{V})\rightarrow V\rightarrow 0,$
@@
Theorem 2.3.
Let $(\Pi,V)$ be a finite length representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$ such that $(\Pi,V)$ extends as a representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. Then each Tate cohomology group $\widehat{H}^{i}(\Gamma,V)$ is a finitely generated representation of $\overline{\mathbb{F}}_{l}[G^{\sigma}(F)]$.
@@
We use induction on the length of $(\Pi,V)$ as an $\overline{\mathbb{F}}_{\ell}[G(F)]$ representation. If $(\Pi,V)$ is an irreducible $G(F)$ representation, then the theorem follows from Lemma 2.2. Let $W$ be an irreducible $G(F)$ subrepresentation of $V$. Let us first consider the case when $W$ is $\Gamma$ stable. The long exact sequence of Tate cohomology associated with the following exact sequence
@@
Since $W$ is irreducible, each $\widehat{H}^{i}(\Gamma,W)$ is a finite length representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$ by Lemma 2.2. Also note that the length of the $G(F)$ representation $V/W$ is strictly less than the length of $V$, inductively we may assume that the Tate cohomology groups $\widehat{H}^{i}(\Gamma,V/W)$ are finite length representations of $G^{\sigma}(F)$. Now, it follows from the above exact sequences that each $\widehat{H}^{i}(\Gamma,V)$ has finite length as a representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$. Let us now consider the case where $W$ is not $\Gamma$ stable. Then
@@
is a $G(F)\rtimes\Gamma$ stable subspace of $V$, where the action of $G(F)$ on the space $W^{\sigma^{i}}$ is twisted by $\sigma^{i}$ for each $i$. Note that $\widehat{H}^{i}(\Gamma,W^{\prime})$ is trivial, for $i\in\{0,1\}$. Now, the long exact sequence of Tate cohomology corresponding to the following short exact sequence of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$ modules
@@
Since the length of $V/W^{\prime}$ is strictly less than that of $V$, the $G^{\sigma}(F)$ representation $\widehat{H}^{i}(\Gamma,V/W^{\prime})$ is of finite length by induction. Thus, we get that the Tate cohomology $\widehat{H}^{i}(\Gamma,V)$ is a finite length representation of $G^{\sigma}(F)$. ∎
2.4.
When the group $G$ is finite, the following lemma gives the trace of $G^{\sigma}$ action on the Tate cohomology of $\sigma$ stable $\ell$-modular representations. This will be useful when we consider the Tate cohomology of a lattice in Shintani lifts.
Lemma 2.4.
Let $G$ be a finite group with an automorphism $\sigma$ of order $\ell$. Let $M$ be a $G\rtimes\langle\sigma\rangle$ stable $\Lambda$-module with $\widehat{H}^{1}(\langle\sigma\rangle,M)=0$. Then
${\rm Tr}(g|\widehat{H}^{0}(\langle\sigma\rangle,M))={\rm Tr}(g|M),$
for all $g\in G^{\sigma}$.
Proof.
Note that the only indecomposable modules of $\mathbb{Z}_{\ell}[\langle\sigma\rangle]$ are given by $\mathbb{Z}_{\ell}$, $\mathbb{Z}_{\ell}[\langle\sigma\rangle]$, and the argumentation ideal $I$. For any $g\in G^{\sigma}$, if $g$ fixes an indecomposable submodule, then the action of $g$ on it has a unique eigenvalue. Thus, ${\rm Tr}(g|M)$ is the sum of ${\rm Tr}(g|W)$, where $W$ is a $g$-stable indecomposable subspace. ∎
3. Finiteness and genericity of Tate cohomology
In this section, we prove some results on the compatibility of Jacquet modules and twisted Jacquet modules with Tate cohomology. Note that these results have been proved in [ND24, Section 6] for the group $G={\rm Res}_{E/F}{\rm GL}_{n}$, where $E/F$ is a finite Galois extension of prime degree $\ell$, and $\sigma$ is an automorphism induced by a generator of the Galois group ${\rm Gal}(E/F)$. But, in our previous work we have not stated them in the generality we need here. The results in this section, although proven over non-Archimedean local fields, are also valid over finite fields. For convenience, we will drop $\Gamma$ in the Tate cohomology groups $\widehat{H}^{i}(\Gamma,M)$ and simply denote it by $\widehat{H}^{i}(M)$.

### v2

@@ 2. Tate cohomology of smooth representations
Let $\Pi$ be a smooth finite length representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. We prove that the Tate cohomology groups $\widehat{H}^{i}(\Gamma,\Pi)$ are finitely generated representations of $G^{\sigma}(F)$. We use Schneider–Stuhler resolutions and their $-\ell$ version by Vignéras for proving this finiteness statement.
@@
Fix an algebraic closure $\overline{\mathbb{Q}}_{\ell}$ of $\mathbb{Q}_{\ell}$ and let $\mathcal{K}$ be the maximal unramified extension of $\mathbb{Q}_{\ell}$ in $\overline{\mathbb{Q}}_{\ell}$. Let $\Lambda$ be the ring of integers of $\mathcal{K}$. All $\Lambda[G(F)]$ modules are assumed to be smooth, i.e., every element is fixed by an open subgroup of $G(F)$. Let $M$ be a $\Lambda[G(F)\rtimes\Gamma]$ module. Let ${\rm Nr}$ be the norm element $\sum_{g\in\Gamma}g\in\Lambda[\Gamma]$. Consider the chain complex of $\Lambda[\Gamma]$ modules
@@
The Tate cohomology groups of $M$, with respect to the action of $\Gamma$, are 222-periodic and we have
@@
The action of $G(F)$ on $M$ induces an action of $G^{\sigma}(F)$ on the Tate cohomology groups $\widehat{H}^{i}(\Gamma,M)$, and we get two $\ell$-modular smooth representations $\widehat{H}^{i}(\Gamma,M)$, for $i\in\{0,1\}$, of the fixed point subgroup $G^{\sigma}(F)$. Similar cohomology groups can be defined when $F$ is a finite field, and we will assume the easy modifications in the definitions.
@@
Let $X$ be a Hausdorff, locally compact, and totally disconnected topological space (an $\ell$-space in the sense of [BZ76, Section 1.1]). Let $G$ be a locally profinite group with an order $\ell$ automorphism $\sigma$, generating a group $\Gamma$ of automorphisms of $G$. Assume that $G\rtimes\Gamma$ acts continuously on $X$. Let $\mathcal{F}$ be a $G\rtimes\Gamma$ equivariant sheaf on $X$. The natural restriction map
@@
is an isomorphism, for $i\in\{0,1\}$ (see [TV16, Proposition 3.3]). Hear, $H^{0}(X,\mathcal{F})$ is the space of global sections of a sheaf $\mathcal{F}$ on $X$. Let $H$ be a closed subgroup of $G$ such that $\sigma(H)=H$. Let $V$ be a $\Lambda[H\rtimes\Gamma]$ module. This data defines a $G\rtimes\Gamma$ equivariant sheaf $\mathcal{F}_{V}$ on $G/H$ whose sections on an open set $U$ of $G/H$ are the sections of the vector bundle $G\times_{H}V\rightarrow G/H$ restricted to $U$. The global sections of the sheaf $\mathcal{F}_{V}$ is the induced representation ${\rm ind}_{H}^{G}V$, and the action of $\Gamma$ is given by
$(\sigma.f)(g)=\sigma\big{(}f(\sigma^{-1}(g))\big{)},$
for all $g\in G$ and $f\in{\rm ind}_{H}^{G}V$. Assume that $(G/H)^{\sigma}=G^{\sigma}/H^{\sigma}$. Then, the restriction to $G^{\sigma}$ map induces a $G^{\sigma}$ equivariant isomorphism
@@
Let $(\Pi,V)$ be a smooth finite length representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. We show that $\widehat{H}^{i}(\Gamma,\Pi)$ is a finitely generated $G^{\sigma}(F)$ representation. We use the $-\ell$ version of Schneider and Stuhler resolutions ([SS97, Theorem II.3.1]) of smooth irreducible representations as discussed in Vignéras ([Vig97, Proposition 2.6]). Let $\mathfrak{X}$ be the extended Bruhat–Tits building of a connected reductive algebraic group $G$ defined over a non-Archimedean local field $F$. Let $d$ be the $F$-rank of $G$. For $0\leq k\leq d$, let $\mathfrak{X}_{k}$ be the set of $k$-dimensional facets of $\mathfrak{X}$, and let $\mathfrak{X}_{(k)}$ be the set of oriented $k$-facets which are tuples of the form $(\mathcal{F},c_{\mathcal{F}})$, where $\mathcal{F}\in\mathfrak{X}_{k}$ and $c_{\mathcal{F}}$ is an orientation of $\mathcal{F}$. Let $e\geq 0$. For a facet $\mathcal{F}$ of $\mathfrak{X}$, Schneider and Stuhler associate a compact open subgroup $U_{\mathcal{F}}^{(e)}$, and let $\underline{V}$ be the coefficient system defined by associating every facet $\mathcal{F}$ to $V^{U^{e}_{\mathcal{F}}}$. Let $C(\mathfrak{X}_{(k)},\underline{V})$ be the set of oriented cellular $k$-chains on $\mathfrak{X}_{(k)}$. Then we have a $G$ equivariant resolution of the representation
$0\rightarrow C(\mathfrak{X}_{(d)},\underline{V})\xrightarrow{\partial_{d-1}}C(\mathfrak{X}_{(d-1)},\underline{V})\xrightarrow{\partial_{d-2}}\cdots\rightarrow C(\mathfrak{X}_{(1)},\underline{V})\xrightarrow{\partial_{0}}C(\mathfrak{X}_{(0)},\underline{V})\xrightarrow{\epsilon}V\rightarrow 0$ (2.2)
for some $e\geq 0$. We refer to [SS97, Section II.2] for details. Note the automorphism $\sigma$ acts on $\mathfrak{X}$ simplicially, i.e., it preserves the sets $\mathfrak{X}_{d}$ and $\mathfrak{X}_{(d)}$.
Lemma 2.3.1.
For $0\leq k\leq d$, the Tate cohomology groups $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(k)},\underline{V}))$ are finitely generated $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$ modules.
@@
Consider the set $\mathfrak{X}_{(k)}$ with discrete topology, and the action of the group $G(F)$ on $\mathfrak{X}_{(k)}$ is continuous. Also note that the sheaf $\underline{V}$ is $G$ and $\Gamma$ equivariant. Since $V$ is an admissible representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$, the stalks of the sheaf $\underline{V}$ are all finite dimensional vector spaces over $\overline{\mathbb{F}}_{\ell}$. Using [TV16, Proposition 3.3], we get that the restriction map
$\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(k)},\underline{V}))\rightarrow C(\mathfrak{X}_{(k)}^{\sigma},\widehat{H}^{i}(\Gamma,\underline{V}))$
is an isomorphism. Let $\mathcal{Y}$ be the Bruhat–Tits building of $G^{\sigma}$, then $\mathfrak{X}^{\sigma}_{k}$ is a subset of facets of $\mathcal{Y}$. Thus, there are only finitely many orbits for the action of $G^{\sigma}$ on $\mathfrak{X}^{\sigma}_{(k)}$. Since the module $\widehat{H}^{i}(\Gamma,V^{U_{F}^{e}})$ is finite, we get that the space $C(\mathfrak{X}_{(k)}^{\sigma},\widehat{H}^{i}(\Gamma,\underline{V}))$ is a finitely generated representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$. This shows the lemma. ∎
In what follows, we need the fact that the submodules of finitely generated smooth $\overline{\mathbb{F}}_{\ell}[H(F)]$ representations are again finitely generated. This is a fundamental result in the theory of smooth complex representations due to Bernstein and in the $-\ell$ case it follows from the deep results of the paper [DHKM24, Corollary 1.3] seen with the work of Dat ([Dat09, Corollary 4.5]).
Lemma 2.3.2.
Let $(\Pi,V)$ be an irreducible smooth representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$. Assume that $(\Pi,V)$ extends as a representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. Then the Tate cohomology groups $\widehat{H}^{i}(\Gamma,V)$ are finitely generated representations of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$.
@@
Note that the resolution (2.2) is $G$ and $\Gamma$ equivariant, and the Tate cohomology groups
@@
are finitely generated representations of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$ by Lemma 2.3.1. For each $2\leq k\leq d$, consider the short exact sequence of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$ modules
$0\rightarrow{\rm Im}(\partial_{k-1})\rightarrow C(\mathfrak{X}_{(k-1)},\underline{V})\rightarrow{\rm Ker}(\partial_{k-2})\rightarrow 0$
Then the corresponding long exact sequence of Tate cohomology groups gives the following hexagonal exact sequence
${\widehat{H}^{0}(\Gamma,{\rm Im}(\partial_{k-1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{0}(\Gamma,C(\mathfrak{X}_{(k-1)},\underline{V}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{0}(\Gamma,{\rm Ker}(\partial_{k-2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{1}(\Gamma,{\rm Ker}(\partial_{k-2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{1}(\Gamma,C(\mathfrak{X}_{(k-1)},\underline{V}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{H}^{1}(\Gamma,{\rm Im}(\partial_{k-1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
which we denote by $S(k)$. Note that ${\rm Im}(\partial_{d-1})$ is identified with $C(\mathfrak{X}_{(d)},\underline{V})$. Since both the Tate cohomology $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d)},\underline{V}))$ and $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d-1)},\underline{V}))$ are finitely generated, it follows from the diagram $S(d)$ that $\widehat{H}^{i}(\Gamma,{\rm Ker}(\partial_{d-2}))$ is also finitely generated. A similar hexagonal diagram of Tate cohomology corresponding to the following short exact sequence of $G(F)\rtimes\Gamma$ modules
$0\rightarrow{\rm Ker}(\partial_{d-2})\rightarrow C(\mathfrak{X}_{(d-1)},\underline{V})\rightarrow{\rm Im}(\partial_{d-2})\rightarrow 0$
shows that $\widehat{H}^{i}(\Gamma,{\rm Im}(\partial_{d-2}))$ is finitely generated. Since $\widehat{H}^{i}(\Gamma,C(\mathfrak{X}_{(d-2)},\underline{V}))$ is finitely generated, it follows from the hexagonal diagram $S(d-1)$ that $\widehat{H}^{i}(\Gamma,{\rm Ker}(\partial_{d-3}))$ is finitely generated. Thus, following a finite number of inductive steps, the hexagonal diagram $S(2)$ shows that the Tate cohomology $\widehat{H}^{i}(\Gamma,{\rm Ker}(\partial_{0}))$ and $\widehat{H}^{i}(\Gamma,{\rm Im}(\partial_{0}))$ are finitely generated. Finally, it follows from the hexagonal diagram of Tate cohomology groups corresponding to the following short exact sequence of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$ modules
$0\rightarrow{\rm Im}(\partial_{0})\rightarrow C(\mathfrak{X}_{(0)},\underline{V})\xrightarrow{\epsilon}V\rightarrow 0,$
@@
Theorem 2.3.3.
Let $(\Pi,V)$ be a finite length representation of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$. Then each Tate cohomology group $\widehat{H}^{i}(\Gamma,V)$ is a finitely generated representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$.
@@
We use induction on the length of $(\Pi,V)$ as an $\overline{\mathbb{F}}_{\ell}[G(F)]$ representation. If $(\Pi,V)$ is an irreducible $G(F)$ representation, then the theorem follows from Lemma 2.3.2. Let $W$ be an irreducible $G(F)$ subrepresentation of $V$. First consider the case when $W$ is $\Gamma$ stable. The long exact sequence of Tate cohomology associated with the following short exact sequence
@@
Since $W$ is irreducible, each Tate cohomology $\widehat{H}^{i}(\Gamma,W)$ is a finitely generated representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$ by Lemma 2.3.2. Also note that the length of the $G(F)$ representation $V/W$ is strictly less than the length of $V$, inductively we may assume that the Tate cohomology groups $\widehat{H}^{i}(\Gamma,V/W)$ are finitely generated representations of $G^{\sigma}(F)$. Now, it follows from the above exact sequences that each $\widehat{H}^{i}(\Gamma,V)$ is a finitely generated representation of $\overline{\mathbb{F}}_{\ell}[G^{\sigma}(F)]$. Let us now consider the case where $W$ is not $\Gamma$ stable. Then
@@
is a $G(F)\rtimes\Gamma$ stable subspace of $V$, where the action of $G(F)$ on the space $W^{\sigma^{i}}$ is twisted by $\sigma^{i}$ for $0\leq i<\ell$. Note that $\widehat{H}^{i}(\Gamma,W^{\prime})$ is trivial, for $i\in\{0,1\}$. Now, the long exact sequence of Tate cohomology groups applied to the following short exact sequence of $\overline{\mathbb{F}}_{\ell}[G(F)\rtimes\Gamma]$ modules
@@
Since the length of $V/W^{\prime}$ is strictly less than that of $V$, the $G^{\sigma}(F)$ representation $\widehat{H}^{i}(\Gamma,V/W^{\prime})$ is finitely generated by induction. Thus, we get that the Tate cohomology group $\widehat{H}^{i}(\Gamma,V)$ is a finitely generated representation of $G^{\sigma}(F)$. ∎
3. Finiteness of Tate cohomology, the Galois automorphisms case
In this section, we prove some results on the compatibility of Jacquet modules and twisted Jacquet modules with Tate cohomology groups. Some of these results were proved in [ND24, Section 6] for the group $G={\rm Res}_{E/F}{\rm GL}_{n}$, where $E/F$ is a Galois extension of prime degree $\ell$, and $\sigma$ is an automorphism induced by a generator of the Galois group ${\rm Gal}(E/F)$. The results in this section, although proved for groups defined over non-Archimedean local fields, are valid for finite groups of Lie type. For convenience, we will drop $\Gamma$ in the Tate cohomology groups $\widehat{H}^{i}(\Gamma,M)$ and simply denote these cohomology spaces by $\widehat{H}^{i}(M)$.